OpenStudy (anonymous):

Can anyone please help me with this problem? \[f(t)\sqrt{t^2+1}\] find the function if it exists

OpenStudy (anonymous):

I suppose the function sqrt(t^2+1) does indeed exist...

OpenStudy (anonymous):

but how do you find the function?

OpenStudy (anonymous):

U appear to have found it, it's right there...

OpenStudy (anonymous):

That is the question

OpenStudy (anonymous):

U are saying find the function but u are telling me what the function is?????

OpenStudy (anonymous):

that is how the problem is written. ???????

OpenStudy (anonymous):

I dont understand what u want me to do, neither does anybody else.

OpenStudy (anonymous):

well i am just writing the problem the way it is written.

OpenStudy (anonymous):

oops find the function value if it exists

OpenStudy (anonymous):

sorry i thought I wrote value .

OpenStudy (anonymous):

The only thing I can say about the function is that it must be positive by definition. Else you have to provide a t to enable calculation of a value for that particular value of t.

OpenStudy (anonymous):

\[f(t)\sqrt{t^2+1} \cdot \sqrt{t^2+1}^{-1} = f(t)\] which is the "value" of the function at t

OpenStudy (anonymous):

it says f(0)= ??????

OpenStudy (anonymous):

That-s better....

OpenStudy (anonymous):

sorry about that estudier.

OpenStudy (anonymous):

So you mean \[ f(t) = \sqrt{t^2+1}\]\[\text{Find $f(0)$}\]

OpenStudy (anonymous):

1

OpenStudy (anonymous):

\[f(0) = \sqrt{(0)^2+1} = \sqrt{1} = 1\]

OpenStudy (anonymous):

Thank you!!!!